# Depth Any Camera: Zero-Shot Metric Depth Estimation from Any Camera

### v1

@@ Abstract
While recent depth foundation models exhibit strong zero-shot generalization, achieving accurate metric depth across diverse camera types—particularly those with large fields of view (FoV) such as fisheye and 360-degree cameras—remains a significant challenge. This paper presents Depth Any Camera (DAC), a powerful zero-shot metric depth estimation framework that extends a perspective-trained model to effectively handle cameras with varying FoVs. The framework is designed to ensure that all existing 3D data can be leveraged, regardless of the specific camera types used in new applications. Remarkably, DAC is trained exclusively on perspective images but generalizes seamlessly to fisheye and 360-degree cameras without the need for specialized training data. DAC employs Equi-Rectangular Projection (ERP) as a unified image representation, enabling consistent processing of images with diverse FoVs. Its key components include a pitch-aware Image-to-ERP conversion for efficient online augmentation in ERP space, a FoV alignment operation to support effective training across a wide range of FoVs, and multi-resolution data augmentation to address resolution disparities between training and testing. DAC achieves state-of-the-art zero-shot metric depth estimation, improving delta-1 accuracy by up to 50% on multiple fisheye and 360-degree datasets compared to prior metric depth foundation models, demonstrating robust generalization across camera types.
@@ 1 Introduction
As illustrated in Fig. 2, large FoV images can be represented in multiple formats, but generating the best-performing undistorted image for perspective-based depth models often leads to substantial FoV loss. Despite these limitations, large FoV inputs are crucial for efficiency-critical downstream applications such as large-scale detection [48, 31], segmentation [54, 51], SLAM [11, 40, 58, 41], interactive 3D scene generation [55], and robotic demonstration capturing [6, 45].
@@
In this paper, we present Depth Any Camera (DAC), a novel zero-shot metric depth estimation framework that enables a depth model trained exclusively on perspective images to generalize across cameras with widely varying FoVs, including fisheye and cameras (see Fig. 1). DAC employs Equi-Rectangular Projection (ERP) as a canonical representation to unify images from diverse FoVs into a shared space. Key innovations include an efficient pitch-aware Image-to-ERP conversion based on grid sampling and Gnomonic Geometry [43], enabling seamless ERP-space data augmentations. Specifically, pitch-aware ERP conversion and pitch-angle augmentation allow perspective data to be projected into high-distortion regions of the ERP space, which are only observable from large-FoV cameras. This enhances the framework’s zero-shot generalization capability, enabling it to extrapolate effectively from the perspective domain to the broader camera domain. To ensure effective learning, we propose a FoV alignment process to adjust diverse-FoV training samples to a predefined ERP patch size, preserving content while minimizing computational overhead. Additionally, multi-resolution augmentation addresses resolution mismatches, allowing the model to learn scale-equivariant features and adapt to a flexible range of testing resolutions.
In summary, our contributions are as follows:
@@
We introduce an efficient perspective-to-ERP patch conversion to enable pitch-aware ERP conversion and ERP-space online augmentation.
We propose a FoV alignment process, facilitating effective training across cameras with diverse FoVs within the unified ERP space and a multi-resolution training strategy to address resolution mismatches between training ERP patches and testing images.
Our method achieves state-of-the-art zero-shot performance on all large FoV testing datasets, delivering up to a 50% improvement in accuracy on multiple indoor fisheye and datasets, showcasing robust generalization across diverse camera types.
@@ 2.2 Depth Estimation from Large FoV Cameras
Depth estimation for fisheye, cameras has grown in popularity, as large FoVs capture richer contextual information that enhances depth estimation [20, 24, 37, 56, 1]. A key challenge for these cameras is managing position-dependent distortions, which vary by camera models. Solutions include deformable CNNs [38, 59, 46], which adapt kernel shapes to handle distortions, and UniFuse [20], which segments ERP images to minimize distortion before merging. More recent methods leverage transformers to handle these distortions [24, 37, 56, 10]. While transformer-based networks have improved in-domain performance, they are approaching saturation, indicating that distortion is not the only challenge. Instead, the lack of large-scale FoV-specific training data is a key bottleneck for generalization. No current methods enable an unified depth estimation model trained on mixed large-scale perspective datasets to achieve zero-shot generalization on ERP or fisheye images.
@@ 3 Notations and Preliminaries
Transformations between standard images and ERP images use Gnomonic Projection transformation [43], which offers a closed-form mapping between normalized image coordinates and spherical coordinates , assuming a tangent plane centered at of a unit sphere. Specifically, as shown in Fig. 5, this mapping is given by:
where is the angular distance of the point from the center of the projection, given by:
@@ 4.1 Pitch-Aware Image-to-ERP Conversion
We propose an efficient approach to directly convert an input image to its corresponding ERP patch through grid sampling combined with gnomonic projection. Assuming an ERP space with height , width , and the image center at latitude , longitude , with a target ERP patch size of , the ERP patch coordinates can be mapped to spherical coordinates as follows:
Using Gnomonic Geometry as presented in Eq. 1 and Eq. 2, we obtain the corresponding normalized image coordinate in the tangent plane touching the unit sphere. To map this coordinate to the actual image coordinate , we apply distortion and projection functions based on the given camera parameters:
where is the distortion function with distortion parameters , and is the projection function with intrinsic parameters . If the input image has no distortion, we can directly apply the projection function to .
@@
One key advantage of this ERP conversion lies in its camera pitch awareness, corresponding to the latitude of the tangent plane center, , as defined in Eq. 1 and Eq. 2. When camera orientation is available or can be estimated [21], perspective data can be projected to various latitudes in the ERP space, enabling the simulation of regions visible only with large field-of-view (FoV) cameras, as shown in Fig. 5, and Supplemental Fig. 7. This pitch-aware conversion is crucial for improving the generalization of trained models to previously unobserved large-FoV data, as demonstrated in Sec 5.3, because neural networks alone have limited capacity to generalize to extrapolated data spaces [47]. For datasets with limited pitch variation [12, 14, 18, 57], pitch augmentation can be efficiently achieved by introducing noise to , generating ERP patches with varying shapes, as shown in Fig. 4 and Fig. 5.
Another notable advantage is the seamless ability to perform online data augmentation efficiently in the ERP space. Common augmentations, such as scaling, rotation, and translation—commonly applied to perspective images—can be directly applied to the normalized image coordinates as follows:
where is a scale factor, is 2D rotation matrix, and is a 2D translation vector corresponding to the applied augmentations. When applying scaling augmentation during training, the ground-truth depth must also be scaled by to maintain scale-consistency in the training data. This is analogous to the image resizing scenario with a fixed camera focal length, as illustrated in Fig. 3.
@@ 4.2 FoV Alignment
To address this challenge, we introduce a simple yet effective FoV Alignment operation that adjusts the FoV of each input image to match the predetermined crop area FoV. Specifically, this process applies a specific scaling augmentation, as described in Sec. 4.1 and Eq. 8, specifically:
where is derived from actual camera parameters, and is ERP patch’s vertical FoV. As illustrated in Fig. 5, this approach allows a single predefined ERP patch size to maximize the inclusion of relevant content and minimize computational waste on background, making it ideal for an efficient training pipeline.
@@ 5.1 Experimental Setup
In-Domain Training Datasets. Table 1 summarizes the datasets used. For indoor experiments, we use Habitat-Matterport 3D (HM3D) [32], Taskonomy [57], and Hypersim [35], totaling 670K perspective images with distinct characteristics. To streamline training, we use the first 50 scenes from HM3D and Taskonomy (tiny versions provided by OmniData [8]). For outdoor data, we use DDAD [14] and LYFT [18], totaling 130K images, with datasets varying widely in FoV, pitch angles, sources, and image quality.
Zero-Shot Testing Datasets. We evaluate DAC on two datasets—Matterport3D [5] and Pano3D-GV2 [2]—and two fisheye datasets—ScanNet++ [52] and KITTI360 [25]—all featuring larger FoVs than perspective images. Our primary experiments focus on these datasets to assess zero-shot generalization to large FoV cameras. Additional evaluations on NYUv2 [27] and KITTI [12] are provided in the supplementary material, demonstrating DAC’s performance on perspective data relative to state-of-the-arts.
@@
Metric3Dv2 [19]: A leading method in zero-shot metric depth estimation, achieving state-of-the-art performance on perspective images through a canonical camera model to standardize datasets.
UniDepth [30]: This recent method generalizes to diverse perspective cameras in zero-shot settings. We test its ability to handle large FoV cameras not included in training.
iDisc [29]: Selected as a baseline due to its use of self-attention and cross-attention modules in a straightforward yet effective network, showing strong in-domain performance. As iDisc alone cannot handle mixed camera parameters, we apply the perspective virtual camera conversion from Metric3Dv2 [19] to train it on varied perspective datasets, then compare it to the same network integrated with DAC for zero-shot generalization on cameras with significantly different FoVs.
@@
For testing resolution, if the original resolution is less than twice the training resolution, we use it directly; for larger resolutions, we maintain the aspect ratio and align it with the training resolution. Based on this rule, we evaluate Matterport3D [5] and Pano3D-GV2 [2] at , ScanNet++ [52] at , and KITTI360 [25] at . For competing methods that are not adaptable to inconsistent resolutions compared to training, we report the higher score obtained from the two resolution settings to ensure fairness.
@@
5.2 Comparison with the State of the Art
In this section, we compare DAC with primary baselines in zero-shot generalization tests on large FoV datasets, with quantitative results reported in Table 2, and qualitative results shown in Fig. 6. In indoor experiments, DAC significantly outperforms pre-trained models UniDepth [30] and Metric3Dv2 [19], even when using a lighter ResNet101 [17] backbone and a much smaller training dataset. DAC achieves superior performance across both datasets and the fisheye dataset ScanNet++ [52]. Compared to the iDisc [29] network trained with the Metric3Dv2 pipeline, DAC shows substantial improvements across all metrics on all datasets. Notably, DAC improves the next-best method by nearly 50% in the most differentiating metric, , on all three datasets.
In outdoor tests, DAC significantly outperforms Metric3Dv2 [53] and UniDepth [30], even with much larger backbones. However, it achieves only marginal improvements over iDisc [29] under the same network configuration, with less pronounced gains compared to indoor settings. This is likely due to the limited camera pitch variance in the outdoor training data (Table 1) and the KITTI360 [25] LiDAR-based ground truth, which primarily covers the central region of a fisheye image (Fig. 6).
@@ 5.3 Ablation Study
Table 3 highlights the pivotal role of pitch-aware ERP conversion in generalizing perspective-trained models to large FoV datasets by projecting inputs to latitude regions unique to such data (Fig. 7). This approach turns the wide pitch angle variance in datasets like HM3D [32] into an advantage. While additional pitch augmentation does not appear essential when the training dataset like HM3D intricately spans a large range of pitch angle. However, its effectiveness varies across datasets, as detailed in Supplemental Table 6.
@@ 7.1 Full Zero-Shot Metric Depth Experiments
Full experiments with a few additional experiments comparing DAC to the state-of-the-arts in zero-short metric depth estimation are shown in Table 5. The additional experiments include:

### v2

@@ Abstract
While recent depth foundation models exhibit strong zero-shot generalization, achieving accurate metric depth across diverse camera types—particularly those with large fields of view (FoV) such as fisheye and 360-degree cameras—remains a significant challenge. This paper presents Depth Any Camera (DAC), a powerful zero-shot metric depth estimation framework that extends a perspective-trained model to effectively handle cameras with varying FoVs. The framework is designed to ensure that all existing 3D data can be leveraged, regardless of the specific camera types used in new applications. Remarkably, DAC is trained exclusively on perspective images but generalizes seamlessly to fisheye and 360-degree cameras without the need for specialized training data. DAC employs Equi-Rectangular Projection (ERP) as a unified image representation, enabling consistent processing of images with diverse FoVs. Its core components include Pitch-aware Image-to-ERP conversion with efficient online augmentation to simulate distorted ERP patches from undistorted inputs, FoV alignment operations to enable effective training across a wide range of FoVs, and multi-resolution data augmentation to further address resolution disparities between training and testing. DAC achieves state-of-the-art zero-shot metric depth estimation, improving accuracy by up to 50% on multiple fisheye and 360-degree datasets compared to prior metric depth foundation models, demonstrating robust generalization across camera types.
@@ 1 Introduction
As illustrated in Fig. 2, large FoV images can be represented in multiple formats, but generating the best-performing undistorted image for perspective-based depth models often leads to substantial FoV loss. Despite these limitations, large FoV inputs are crucial for efficiency-critical downstream applications such as large-scale detection [52, 34], segmentation [58, 55], SLAM [11, 44, 62, 45], interactive 3D scene generation [59], and robotic demonstration capturing [6, 49, 14].
@@
In this paper, we present Depth Any Camera (DAC), a novel zero-shot metric depth estimation framework that enables a depth model trained exclusively on perspective images to generalize across cameras with widely varying FoVs, including fisheye and cameras (see Fig. 1). DAC employs Equi-Rectangular Projection (ERP) as a canonical representation to unify images from diverse FoVs into a shared space. A key innovation is the introduction of an efficient Pitch-aware Image-to-ERP conversion based on grid sampling and Gnomonic Geometry [47], enabling seamless ERP-space data augmentations. Specifically, pitch-aware ERP conversion with pitch-angle augmentation projects perspective images into high-distortion regions of the ERP space, effectively simulating observations unique to large-FoV cameras. This enhances DAC’s zero-shot generalization, allowing it to extrapolate beyond the perspective domain to a broader range of camera types. To facilitate learning from mixed datasets, we propose a FoV alignment process that normalizes diverse-FoV training samples to a predefined ERP patch size, preserving content while minimizing computational overhead. Additionally, multi-resolution augmentation is applied to address resolution mismatches, allowing the model to learn scale-equivariant features and adapt to a flexible range of testing resolutions. In summary, our contributions are as follows:
@@
We introduce an efficient pitch-aware Image-to-ERP conversion that simulates the high-distortion characteristics of large-FoV cameras from perspective inputs, enhancing zero-shot generalization.
We develop a FoV alignment process that enables effective training across cameras with diverse FoVs within a unified ERP space, along with a multi-resolution training strategy to address resolution mismatches between training ERP patches and testing images.
Our method achieves State-of-The-Art (SoTA) zero-shot performance on all large FoV testing datasets, delivering up to a 50% improvement in accuracy on indoor fisheye and datasets, showcasing strong generalization across diverse camera types.
@@ 2.2 Depth Estimation from Large FoV Cameras
Depth estimation for fisheye, cameras has grown in popularity, as large FoVs capture richer contextual information that enhances depth estimation [21, 26, 41, 60, 1]. A key challenge for these cameras is managing position-dependent distortions, which vary by camera models. Approaches to address this include deformable CNNs [42, 63, 50], which adapt kernel shapes to compensate for distortions, as well as methods that segment ERP images to reduce distortion effects before merging [21, 38]. More recent methods leverage transformers to handle these distortions [26, 41, 60, 10]. While transformer-based networks have improved in-domain performance, they are approaching saturation, indicating that distortion is not the only challenge. Instead, the lack of large-scale FoV-specific training data is a key bottleneck for generalization. No current methods enable an unified depth estimation model trained on mixed large-scale perspective datasets to achieve zero-shot generalization on ERP or fisheye images.
@@ 3 Notations and Preliminaries
Transformations between standard images and ERP images use Gnomonic Projection transformation [47], which offers a closed-form mapping between tangent image coordinates and spherical coordinates , assuming a tangent plane centered at of a unit sphere. Specifically, as shown in Fig. 5, this mapping is given by:
where represents a point on the unit sphere, and is the angular distance between and , can be calculated as:
@@ 4.1 Pitch-Aware Image-to-ERP Conversion
An input image can be efficiently converted to its corresponding ERP patch through grid sampling combined with gnomonic projection. Assuming an ERP space with height , width , and the image center at latitude , longitude , with a target ERP patch size of , the ERP patch coordinates can be mapped to spherical coordinates as , and . Using Gnomonic Geometry presented in Eq. 1 and Eq. 2, we obtain the corresponding normalized image coordinate in the tangent plane and on the unit sphere. To map this coordinate to the actual image coordinate , we apply distortion and projection functions based on the given camera parameters:
where is the distortion function with distortion parameters , and is the projection function with intrinsic parameters . If the input image has no distortion, projection function is directly applied to . Details on applying distortion models are included in Supplemental Sec. 8.
@@
This ERP conversion enables a powerful training pipeline when the latitude of the tangent plane center is defined by the camera’s pitch angle in training. When camera orientation is available or can be estimated [22], perspective data can be projected to various latitudes in the ERP space, enabling the simulation of regions uniquely visible from large-FoV cameras, as shown in Fig. 5, and Supplemental Fig. 7. This pitch-aware conversion is crucial for improving the generalization of trained models to previously unobserved large-FoV data, as demonstrated in Sec. 5.3, because neural networks alone have limited capacity to generalize to extrapolated data spaces [51].
Another notable advantage is the seamless ability to perform online augmentation efficiently in the ERP space. For datasets with limited pitch variation [12, 15, 19, 61], a unique pitch augmentation can be efficiently applied by adding noise to , generating ERP patches with varying shapes, as shown in Fig. 4. In addition, common augmentations, such as scaling, rotation, and translation—commonly applied to perspective images—can be directly applied to the normalized image coordinates as follows:
where is a scale factor, is 2D rotation matrix, and is a 2D translation vector corresponding to the applied augmentations.
@@ 4.2 FoV Alignment
To address this challenge, we introduce a simple yet effective FoV Alignment operation that adjusts the FoV of each input image to match the predetermined crop area FoV. Specifically, this process applies a specific scaling augmentation, as described in Sec. 4.1 and Eq. 6, specifically , where is derived from actual camera parameters, and is ERP patch’s vertical FoV. As illustrated in Fig. 5, this approach allows a single predefined ERP patch size to maximize the inclusion of relevant content and minimize computational waste on background, making it ideal for an efficient training pipeline.
@@ 5.1 Experimental Setup
In-Domain Training Datasets. For indoor experiments, we use Habitat-Matterport 3D (HM3D) [35], Taskonomy [61], and Hypersim [39], totaling 670K perspective images with distinct characteristics. To streamline training, we use the first 50 scenes from HM3D and Taskonomy (tiny versions provided by OmniData [8]). For outdoor data, we use DDAD [15] and LYFT [19], totaling 130K images. Table 1 summarizes the datasets in use, showing varying distributions in FoV, pitch angles, sources, and image quality.
Zero-Shot Testing Datasets. We evaluate DAC on two datasets—Matterport3D [5] and Pano3D-GV2 [2]—and two fisheye datasets—ScanNet++ [56] and KITTI360 [27]—all featuring larger FoVs than perspective images. Our primary experiments focus on these datasets to assess zero-shot generalization to large FoV cameras. Additional evaluations on NYUv2 [30] and KITTI [12] are provided in the supplementary material, demonstrating DAC’s performance on perspective data relative to SoTA methods.
@@
Metric3Dv2 [20]: A SoTA foundation model in zero-shot metric depth estimation, built upon perspective canonical camera model to standardize datasets.
UniDepth [33]: A more recent SoTA foundation depth model leveraging network designs to handle diverse camera parameters. We test its ability to handle large FoV cameras not included in training.
iDisc [32]: Selected as a network baseline due to its use of self-attention and cross-attention modules in a straightforward yet effective network, and strong in-domain performance. As iDisc alone does not handle mixed camera parameters, we train it with Metric3Dv2 [20], and compare it to the same network trained with ours.
@@
For testing resolution, if the original resolution is less than twice the training resolution, we use it directly; for larger resolutions, we maintain the aspect ratio and align it with the training resolution. Based on this rule, we evaluate Matterport3D [5] and Pano3D-GV2 [2] at , ScanNet++ [56] at , and KITTI360 [27] at . For competing methods that are not adaptable to inconsistent resolutions compared to training, we report the higher score obtained from the two settings to ensure fairness.
@@
5.2 Comparison with the SoTA
In this section, we compare DAC with primary baselines in zero-shot generalization tests on large FoV datasets, with quantitative results reported in Table 2, and qualitative results shown in Fig. 6. In indoor experiments, DAC significantly outperforms pre-trained models UniDepth [33] and Metric3Dv2 [20], even when using a lighter ResNet101 [18] backbone and a much smaller training dataset. DAC achieves superior performance across both datasets and the fisheye dataset ScanNet++ [56]. Compared to the iDisc [32] network trained with the Metric3Dv2 pipeline, DAC shows substantial improvements across all metrics on all datasets. Notably, DAC improves the next-best method by nearly 50% in the most differentiating metric, .
In outdoor tests, DAC significantly outperforms Metric3Dv2 [57] and UniDepth [33], even with much larger backbones. However, it achieves only marginal improvements over iDisc [32] under the same network configuration, with less pronounced gains compared to indoor settings. This is likely due to the limited camera pitch variance in the outdoor training data (Table 1), reducing the ability to simulate highly distorted regions. Moreover, KITTI360 [27] LiDAR points are concentrated in less distorted areas (Fig. 6), making the evaluation less distinctive.
@@ 5.3 Ablation Study
Table 3 highlights the pivotal role of pitch-aware ERP conversion in generalizing perspective-trained models to large FoV datasets by effectively simulating high-distortion regions uniquely observed in large FoV images (Fig. 7). This approach turns the wide pitch angle variance in datasets like HM3D [35] into an advantage. While additional pitch augmentation does not appear essential when the training dataset like HM3D intricately spans a large range of pitch angle. However, its effectiveness varies across datasets, as detailed in Supplemental Table 6.
@@ 7.1 Full Zero-Shot Metric Depth Experiments
Full experiments with a few additional experiments comparing DAC to the SoTA methods in zero-short metric depth estimation are shown in Table 5. The additional experiments include:
@@ 7.4 Zero-Shot Test of Perspective Depth Model on Distorted Images
8 On Applying Camera Distortion Models
As described in Sec. 4.1, the conversion between actual image and the ERP can seamlessly handle different distortion models. In this section, we illustrate how we apply to two typical fisheye models: KB (OpenCV Fisheye) model [23] and MEI model [29].
8.1 KB Model
KB model typically includes distortion parameters . Applying KB model to our Eq. 4 can start from mapping our definition in Eq. 1 and Eq. 2 to the original KB model notations to get:
However, the direct use of can face numerical issue when the FOV is near , when the dividing of approaches 0 in computing them. A more numerical stable version supporting KB at is to use the numerators in Eq. 1 and Eq. 2, denoted as . Then we can rewrite:
where we can keep the ratios consistent between two approaches, while avoiding numeric issues caused by dividing .
The remaining process is exactly the same as the original KB model. Fisheye distortion is applied as:
The distorted point coordinates are where
Finally, given a intrinsic model including as parameters, the conversion into pixel coordinates can be written as:
8.2 MEI Model
MEI model is general more complex by including parameters , where an additional shift parameter is applied so that the model handle even larger FOV camera, and are including tangential distortion.
Mapping our definitions to MEI model is even simpler. Note that actually describe a point lying on the unit sphere, equalizing the Cartesian coordinates converted from the spherical coordinates. The projection coordinates are computed as:
The distortion is then applied as:
Tangential distortion is further applied as:
The later projection is applied the same way as KB model.